OpenStudy (anonymous):

When you graph a parabola with a vertical directrix and the directrix is x=-1/12y^2 and the form they give you is x=a(y-k)^2+h how do write x=-1/12y^2 to that format

hartnn (hartnn):

you need the equation of parabola ?

OpenStudy (anonymous):

yea like i know the format i just dont know what numbers to plug in my original equation is x=-1/12y^2

OpenStudy (anonymous):

would it look like this x=3(y-0)+_?

hartnn (hartnn):

it is in that format only, x= (-1/12) (y-0)^2 +0

hartnn (hartnn):

so, here a = -1/12 k=0 h=0

OpenStudy (anonymous):

oh ok i understand now thanks

hartnn (hartnn):

you're welcome ^_^